OpenStudy (gokuporter):

What is the measure of an exterior angle of a regular 7-sided polygon? Round to the nearest tenth of a degree.

OpenStudy (gokuporter):

40.0

OpenStudy (gokuporter):

I know the measure is 360

OpenStudy (vuriffy):

Isn't there a formula for this like: 180(n-2) find the sum divide by the original sides then subtract 180 and the other value you receive from dividing by the original sides.

OpenStudy (vuriffy):

180(7-2) 5*180 = 900 900 / 7 = 128.57 180 - 128.6 = 51.4 I am be incorrect, it's been awhile since I have gone over this.

OpenStudy (vuriffy):

I may be incorrect*

OpenStudy (gokuporter):

I thought that formula was for interior angles

OpenStudy (vuriffy):

Well, let's see: 51.4 * 7 = 359.8.

OpenStudy (vuriffy):

Pretty close to 360. The formula I used may be wrong.

OpenStudy (vuriffy):

Ahh, yes, I believe that is for interior angles.

OpenStudy (gokuporter):

I think you are correct since I trust you, I will inform you if your answer is right.

OpenStudy (vuriffy):

Give me a second.

OpenStudy (vuriffy):

Woah... crazy enough; it says 360/n (n equals the number of angles) 360/7 = 51.4

OpenStudy (vuriffy):

This may just be a coincidence.

OpenStudy (gokuporter):

Lets see

OpenStudy (gokuporter):

You were correct :)

OpenStudy (vuriffy):

Okay, everything I have discovered says to divide 360 by the number of sides, or angles which still works in this case of 360/7 = 51.4. I believe what I first stated is for interior angles, but both have the same value I believe for this polygon!

OpenStudy (gokuporter):

I got the other question wrong

OpenStudy (vuriffy):

Which?

OpenStudy (gokuporter):

http://openstudy.com/study#/updates/57e1eadbe4b0dc03c97462fe This one